OpenStudy (anonymous):

Match each set of ordered pairs with their rise over run. Match Term Definition negative three comma four and negative two comma one A) negative three three comma negative four and two comma one B) negative one negative three comma four and two comma one C) negative three−fifths three comma negative four and negative two comma one D) negative five

Directrix (directrix):

@nyprincess2001 Would you rewrite these for easier reading, please? By that, I mean, write: "negative three comma four and negative two comma one" as (-3,4) and (-2,1). Thanks.

OpenStudy (anonymous):

(-3,4), (-2,1) A.) -3 (3,-4), (2,1) B.) -1 (-3,4), (2,1) C.) -3/5 (3,-4), (-2,1) D.) -5

Directrix (directrix):

Let's get the rise over run (slope) of the first pair of points: (-3,4), (-2,1) Rise is the difference of the y coordinates. (4 - 1) = 3 Run is the difference of the x coordinates. (-3 -(-2)) = (-3 + 2) = -1 For the points (-3,4), (-2,1) the rise over run is 3/(-1) = -3. Where does that match? @nyprincess2001

Directrix (directrix):

@nyprincess2001 I'm waiting for your reply. Be brave; if what you write is wrong, you can always delete it.

OpenStudy (anonymous):

i THINK IT IS... A or B

Directrix (directrix):

From the work I did above (check it for accuracy) this happened: For the points (-3,4), (-2,1) the rise over run is 3/(-1) = -3. So that option is A.) -3 Okay, @nyprincess2001 it is your turn to get the rise over run for the next pair of points: (3,-4), (2,1) ---> Yours

Directrix (directrix):

Post what you get and we'll compare answers, okay?

OpenStudy (anonymous):

I got b

Directrix (directrix):

I did not get B @nyprincess2001 Look at my work.